# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 1.5 Technical Overview
We improve the construction, eliminating the double-logarithmic factors in the seed length, by replacing the last INW generator with recursive weighted pseudorandom reductions. We start from the reduced ROBP of length $n_{1}$ and alphabet $\{0,1\}^{s_{1}}$, where $n_{1}=m$ and $s_{1}=d$. We set $\varepsilon^{\prime}=(\varepsilon/(2K))^{2}$ as the target error for upcoming weighted reductions. At the $i$th level of the recursion, assume the current permutation ROBP has length $n_{i}$ and alphabet $\{0,1\}^{s_{i}}$. We use the above WPRG $(G,w)$ to fool it except that we set the target error to $\varepsilon^{\prime}$ and the moderate error to $\tau=2^{\frac{C\log(1/\varepsilon^{\prime})}{\log^{2}n_{i}}}$, where $C>0$ is a constant. Denote $(G,w)$ with these parameters as $\mathsf{R}^{(i)}$. Then we can view $\mathsf{R}^{(i)}$ as a weighted pseudorandom reduction. It can reduce the current permutation ROBP to permutation ROBPs of length $n_{i+1}$ and alphabet $\{0,1\}^{s_{i+1}}$. Here $n_{i+1}=O\left(\log n_{i}\frac{\log(1/\varepsilon^{\prime})}{\log(1/\tau)}\right)=\log^{3}n_{i}$ if we set $C$ in $\tau$ to be large enough. And $s_{i+1}$ is identical to the seed length of the INW generator $\mathsf{PRG}(\cdot)$ used in $\mathsf{R}^{(i)}$ , thus $s_{i+1}=s_{i}+O(\log n_{i}\cdot(\log\log n_{i}+\log(1/\tau)))=s_{i}+O\left(\frac{\log(1/\varepsilon^{\prime})}{\log n_{i}}\right)$. We emphasize that the increment from $s_{i}$ to $s_{i+1}$ is not significant, hence we do not need an alphabet reduction to control the alphabet size. This level of recursion has weight $W_{i}=2^{O(\log^{3}n_{i})}$ and costs $\log W_{i}$ bits of randomness. We do the recursion for $l-1$ times, where $l$ is chosen in a way such that $n_{l}=O(1)$. The total weight of the recursion is bound by $2^{O(\log^{3}n_{1})}$ since $n_{i}$ decreases quickly. Notice that this is negligible compared to $\varepsilon/K$, which means that $\varepsilon^{\prime}=(\varepsilon/(2K))^{2}$ is small enough to deduce an $\varepsilon$-WPRG. Finally, we calculate the random bits used. The overall random bits used in our WPRG have three parts, the bits for the final ROBP of length $n_{l}$ and alphabet $\{0,1\}^{s_{l}}$, the bits used in the recursion, and the bits used in the first weighted reduction to choose an addend of the matrix iteration polynomial. For the first part, we have $s_{l}=s_{1}+O(\log(1/\varepsilon^{\prime}))\cdot\sum_{i=1}^{l-1}\frac{1}{\log n_{i}}=s_{1}+O(\log(1/\varepsilon^{\prime}))$. For the second part, one can deduce that the number of bits is $\leq O(\log^{3}n_{1})$, since $\forall i,n_{i+1}=\log^{3}n_{i}$. This is negligible compared to $\log(1/\varepsilon^{\prime})$. The third part requires $\log K$ bits. Therefore, the overall random bits used in our WPRG is $s_{l}\cdot n_{l}+O(\log^{3}n_{1})+\log W=O(s_{1}+\log(1/\varepsilon^{\prime})+\log K)=O(s+\log n\cdot(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon))$.
Derandomization of short-wide regular ROBPs
Our derandomization adapts our WPRG for permutation ROBPs and utilizes the rotation technique which is introduced in [RVW00, RV05] and commonly used in [AKM+20, CHL+23, CL24]. Recall that the rotation technique in [AKM+20, CHL+23, CL24] is used to replace the expander walks in INW and also those one-step transitions on the input ROBP in a white-box way, such that one can use this adapted WPRG to fool the regular ROBP as if one is using the original WPRG to fool a permutation ROBP. Although our WPRG is constructed by recursive weighted pseudorandom reductions, notice that for each level of the recursion, it still inherits the expander walk structure of INW. We show that one can use a sequence of different rotations for different levels of the recursion and different positions of the ROBP to adapt the WPRG such that it can simulate random walks on the input regular ROBP. The parameters basically remain the same as that of our WPRG for permutation ROBPs.
@@ 2.1 Some notations
For convenience of description, we denote $\mathcal{B}_{(n,s,w)}$ as the class of ROBPs with length $n$, alphabet size $2^{s}$ and width $w$. We denote $\mathcal{P}_{(n,s)}$ as the class of permutation ROBPs with unbounded width and only one accept node, where $n$ is the length $s$ is the log-size of the alphabet. We denote $\mathcal{R}_{(n,s,w)}$ as the class of regular ROBPs with length $n$, alphabet size $2^{s}$ and width $w$.
@@ 2.2 Weighted pseudorandom reductions
Here we recall the definition of Weighted Pseudorandom Reduction.
@@
Let $\mathcal{F}_{0}$ be a class of functions ($\left(\{0,1\}^{s_{0}}\right)^{n_{0}}\rightarrow\mathbb{R}$. Let $\mathcal{F}_{simp}$ be a class of functions ($\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\rightarrow\mathbb{R}$. A $(d,K,\varepsilon)$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{simp}$ is a tuple $(\mathsf{R},w)$, in which $\mathsf{R}:\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\times\{0,1\}^{d}\rightarrow\left(\{0,1\}^{s_{0}}\right)^{n_{0}}$ and $w:\{0,1\}^{d}\rightarrow\mathbb{R}$. Furthermore, for every $f\in\mathcal{F}_{0}$, we have:
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d}}\sum_{i\in\{0,1\}^{d}}w(i)\mathbb{E}_{U^{\prime}}[f(\mathsf{R}(U^{\prime},i))]\right|\leq\varepsilon,$
$\forall i,|w(i)|\leq K,$
@@
Let $\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{2}$ be classes of boolean functions within $\{0,1\}^{s_{0}}\to\mathbb{R},\{0,1\}^{s_{1}}\to\mathbb{R},\{0,1\}^{s_{2}}\to\mathbb{R}$ respectively. Let $(\mathsf{R}^{(1)},w^{(1)})$ be an explicit $(d_{1},K_{1},\varepsilon_{1})$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{1}$ and $(\mathsf{R}^{(2)},w^{(2)})$ be an explicit $(d_{2},K_{2},\varepsilon_{2})$-weighted pseudorandom reduction from $\mathcal{F}_{1}$ to $\mathcal{F}_{2}$. Then the composition $(\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)},w^{(1)}\cdot w^{(2)}):$
@@
$w^{(1)}\cdot w^{(2)}(i_{1},i_{2}):=w^{(1)}(i_{1})\cdot w^{(2)}(i_{2}),$
@@
Let $\mathsf{R}=\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)}$ and $w=w^{(1)}\cdot w^{(2)}$. For any $f\in\mathcal{F}_{0}$, we have:
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d_{1}+d_{2}}}\sum_{i_{1}\in\{0,1\}^{d_{1}},i_{2}\in\{0,1\}^{d_{2}}}w(i_{1},i_{2})\mathbb{E}_{U_{s_{2}}}\left[f(\mathsf{R}_{(i_{1},i_{2})}(U_{s_{2}}))\right]\right|$
$\leq\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d_{1}}}\sum_{i_{1}\in\{0,1\}^{d_{1}}}w^{(1)}(i_{1})\mathbb{E}_{U_{s_{1}}}\left[f(\mathsf{R}^{(1)}_{i_{1}}(U_{s_{1}}))\right]\right|$
$+\frac{1}{2^{d_{1}}}\sum_{i_{1}\in\{0,1\}^{d_{1}}}|w^{(1)}(i_{1})|\left|\mathbb{E}_{U_{s_{1}}}\left[f(\mathsf{R}^{(1)}_{i_{1}}(U_{s_{1}}))\right]-\frac{1}{2^{d_{2}}}\sum_{i_{2}\in\{0,1\}^{d_{2}}}w^{(2)}(i_{2})\mathbb{E}_{U_{s_{2}}}\left[f(\mathsf{R}^{(2)}_{i_{2}}(U_{s_{2}}))\right]\right|$
@@
Furthermore, $|w(i_{1},i_{2})|\leq|w^{(1)}(i_{1})|\cdot|w^{(2)}(i_{2})|\leq K_{1}K_{2}$. For any $i_{1},i_{2}$, $x\mapsto f(\mathsf{R}^{(1)}_{i_{1}}(x))\in\mathcal{F}_{1}$, sp we can apply $\mathsf{R}^{(2)}_{i_{2}}$ to this mapping, therefore $x\mapsto f(\mathsf{R}_{(i_{1},i_{2})}(x))\in\mathcal{F}_{2}$. The seed length is $d_{1}+d_{2}$. The function is explicit since both components are explicit. ∎
@@
For any positive integer $k$, let $\mathcal{F}_{0},\mathcal{F}_{1},\ldots,\mathcal{F}_{k}$ be classes of boolean functions within $\{0,1\}^{s_{0}}\to\mathbb{R},\{0,1\}^{s_{1}}\to\mathbb{R},\ldots,\{0,1\}^{s_{k}}\to\mathbb{R}$ respectively. Let $(\mathsf{R}^{(1)},w^{(1)}),\ldots,(\mathsf{R}^{(k)},w^{(k)})$ be explicit $(d_{1},K_{1},\varepsilon_{1}),\ldots,(d_{k},K_{k},\varepsilon_{k})$-weighted pseudorandom reductions from $\mathcal{F}_{0}$ to $\mathcal{F}_{1},\ldots,\mathcal{F}_{k}$ respectively. Then the composition $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k)},w^{(1)}\cdot\ldots\cdot w^{(k)}):$
@@
$w^{(1)}\cdot\ldots\cdot w^{(k)}(i_{1},\ldots,i_{k})=w^{(1)}(i_{1})\cdot\ldots\cdot w^{(k)}(i_{k}),$
@@
We prove this by induction on $k$. The base case $k=2$ is shown by Lemma 2.2. Suppose the statement holds for $k-1$, then $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k-1)},w^{(1)}\cdot\ldots\cdot w^{(k-1)})$ is an explicit $(\sum_{i=1}^{k-1}d_{i},\prod_{i=1}^{k-1}K_{i},\sum_{i=1}^{k-1}\left(\prod_{j=1}^{i-1}K_{j}\right)\varepsilon_{i})$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{k-1}$. Then we can apply Lemma 2.2 on $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k-1)},w^{(1)}\cdot\ldots\cdot w^{(k-1)})$ and $(\mathsf{R}^{(k)},w^{(k)})$, which gives the desired result. ∎

### v3

@@ 1.5 Technical Overview
We emphasize that the alphabet bit-length increment from $s_{i}$ to $s_{i+1}$ is not significant. In fact $s_{l}=s_{1}+O(\log(1/\varepsilon^{\prime}))\cdot\sum_{i=1}^{l-1}\frac{1}{\log n_{i}}=s_{1}+O(\log(1/\varepsilon^{\prime}))$. Also notice that the overall weight $K_{1}K_{2}K_{3}\ldots K_{l}$. Here $K_{2}K_{3}\ldots K_{l}$ is bound by $2^{O(\log^{3}n_{1})}$ since $n_{i}$ decreases quickly. This is negligible compared to $\varepsilon/K_{1}$, which means that $\varepsilon^{\prime}=(\varepsilon/(2K_{1}))^{2}$ is small enough to deduce an $\varepsilon$-WPRG. Finally, we calculate the seed length. The overall random bits used in our WPRG have two parts, the bits for the final ROBP of length $n_{l}$ and alphabet $\{0,1\}^{s_{l}}$ and the randomness used in the weighted reductions. For the first part, we need $s_{l}\cdot n_{l}$ bits. For the second part, we need $\sum_{i}d_{i}=\log K_{1}+\sum_{i=2}^{l}\log K_{i}=\log K_{1}+O(\log^{3}{n_{1}})$ bits. Therefore, one can deduce the overall random bits used in our WPRG is $s_{l}\cdot n_{l}+\sum_{i}d_{i}=O(s+\log n\cdot(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon))$.
Derandomization for short-wide regular ROBPs
To attain our derandomization, notice that if we only want to derandomize permutation ROBPs with multiple accept nodes then we can directly use our WPRG for such permutation ROBPs. Recall that the seed length is $O\left(s+\log n\left(\log\log n+\sqrt{\log(w/\varepsilon)}\right)+\log(w/\varepsilon)\right)$. Hence it is $O(s+\log w)$ if $n=2^{O(\sqrt{\log w})}$, $\varepsilon=1/\operatorname{\mathrm{poly}}w$. In fact, for regular ROBPs, we essentially do the same thing, except that we apply some white-box techniques such that the framework of our WPRG for permutation ROBPs can also work for regular ROBPs.
To derandomize regular ROBPs with binary alphabet, we show that there is a logspace algorithm that can transform a regular ROBP into a permutation ROBP which has the same expectation (though it probably computes a different function). Then we approximation this expectation with our WPRG for permutation ROBPs.
The hard case is to derandomize regular ROBPs with larger alphabets. Let $f$ be a regular ROBP with length $n$, width $w$, alphabet bit-length $s$. We apply $l$ reductions as the following. For every $p\in[l]$, let $f_{i_{1},\dots,i_{p-1}}$ be a reduced regular ROBP of length $n_{p-1}$, with its stochastic matrices ${{\mathbf{M}}_{l\ldots r}}$ from its layer $l$ to layer $r$, $\forall l,r\in[n_{p}]$. $f_{i_{1},\dots,i_{p-1}}$ is $f$ if $p-1=0$. For every $0\leq l<r\leq n_{p-1}$ one can construct a $\tau_{p}$-sv-approximation $\widetilde{{\mathbf{M}}}_{l\ldots r}$ of ${{\mathbf{M}}_{l\ldots r}}$ by the derandomizing square method of [AKM+20, HPV21a, CHL+23], if $f_{i_{1},\dots,i_{p-1}}$ has a two-way labeling. A crucial point is that $\widetilde{{\mathbf{M}}}_{l\ldots r}$ corresponds to the stochastic matrix of a regular bipartite graph $\widetilde{G}_{l\ldots r}$ also with a two-way labeling, as long as $f_{i_{1},\dots,i_{p-1}}$ has a two-way labeling. Recall that a two-way labeling for an ROBP, defined by [RVW00, RV05], requires that every edge has two labels, each for one direction, such that for every vertex all its out-going labels are distinct, and all its in-coming labels are also distinct. And a two-way labeling naturally supports a rotation technique which, whenever the pseudorandom walk arrives at a new node $v$ from $u$ through an edge $e$, switches the label of $e$ from the out-going label of $u$ to the incoming label of $v$. In this way, though the program is only regular, it can still be fooled by a pseudorandom walk from [AKM+20, CHL+23] with such rotations. Using an error reduction polynomial of [CHL+23], one can decompose ${{\mathbf{M}}_{0\ldots n_{p-1}}}$ into a signed sum:
${{\mathbf{M}}_{0\ldots n_{p-1}}}\approx\sum_{i_{p}\in[K_{p}]}\sigma^{(p)}_{i_{p}}\cdot\prod_{t=1}^{n_{p}}\widetilde{{\mathbf{M}}}_{m^{(p)}_{i_{p},t-1}\ldots m^{(p)}_{i_{p},t}},$
where $0=m^{(p)}_{i_{p},0}<m^{(p)}_{i_{p},1}<\cdots<m^{(p)}_{i_{p},n_{p}}=n_{p-1}$ are indices partitioning $f_{i_{1},\dots,i_{p-1}}$ into $n_{p}$ segments, $K_{p}$ denotes the number of terms, and $\sigma^{(p)}_{i_{p}}\in\{-1,0,1\}$ are signs. Each product in the sum corresponds to a concatenation of regular bigraphs such that the product can also be viewed as a regular ROBP $f_{i_{1},\dots,i_{p}}$ of length $n_{p}$:
$f_{i_{1},\dots,i_{p}}=\widetilde{G}_{m^{(p)}_{i_{p},0}\ldots m^{(p)}_{i_{p},1}}\circ\widetilde{G}_{m^{(p)}_{i_{p},1}\ldots m^{(p)}_{i_{p},2}}\circ\cdots\circ\widetilde{G}_{m^{(p)}_{i_{p},n_{p}-1}\ldots m^{(p)}_{i_{p},n_{p}}}.$
$f_{i_{1},\dots,i_{p}}$ again has a two-way labeling since it is the concatenation of bigraphs with two-way labelings. If consider all the reductions, then finally we can approximate $\mathbb{E}[f]$ with a signed sum of expectations of those reduced ROBPs:
$\mathbb{E}[f]\approx\sum_{i_{1}\in[K_{1}],\dots,i_{l}\in[K_{l}]}\sigma^{(1)}_{i_{1}}\cdots\sigma^{(l)}_{i_{l}}\cdot\mathbb{E}[f_{i_{1},\dots,i_{l}}],$
where $\forall p\in[l],\forall i_{p},\sigma^{(p)}_{i_{p}}$ is a corresponding sign from the $p$-th reduction. It turns out all relevant parameters (e.g. the length and alphabet of the reduced ROBPs, the signs, the seed length) can be essentially the same as those parameters of our WPRG for permutation ROBPs. Also one can see that since $f$ has a two-way labeling then $\forall p\in[l]$, every reduced program $f_{i_{1},\dots,i_{p}}$ has a two-way labeling by the property mentioned above. So finally for each $f_{i_{1},\dots,i_{l}}$, whose length is $n_{l}=O(1)$, one can apply a true random walk using rotations supported by the two-way labeling, and the expectation can be approximated by a weighted sum of the results of these random walks. By a more detailed analysis, we show that these rotations are again space efficient so that the overall space complexity is only linear of the seed length.
@@ 2.1 Some notations
For convenience of description, we denote $\mathcal{B}_{(n,s,w)}$ as the class of ROBPs with length $n$, alphabet size $2^{s}$ and width $w$. We denote $\mathcal{P}_{(n,s)}$ as the class of permutation ROBPs with unbounded width and only one accept node, where $n$ is the length $s$ is the bit-length of the alphabet. We denote $\mathcal{R}_{(n,s,w)}$ as the class of regular ROBPs with length $n$, alphabet size $2^{s}$ and width $w$.
@@ 2.2 Weighted pseudorandom reductions
Here we state the definition of weighted pseudorandom reduction.
@@
Let $\mathcal{F}_{0}$ be a class of functions ($\left(\{0,1\}^{s_{0}}\right)^{n_{0}}\rightarrow\mathbb{R}$. Let $\mathcal{F}_{simp}$ be a class of functions ($\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\rightarrow\mathbb{R}$. A $(d,K,\varepsilon)$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{simp}$ is a tuple $(\mathsf{R},w)$, in which $\mathsf{R}:\left(\{0,1\}^{s_{1}}\right)^{n_{1}}\times\{0,1\}^{d}\rightarrow\left(\{0,1\}^{s_{0}}\right)^{n_{0}}$ and $\sigma:\{0,1\}^{d}\rightarrow\mathbb{R}$. Furthermore, for every $f\in\mathcal{F}_{0}$, we have:
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d}}\sum_{i\in\{0,1\}^{d}}\sigma(i)\mathbb{E}_{U^{\prime}}[f(\mathsf{R}(U^{\prime},i))]\right|\leq\varepsilon,$
$\forall i,|\sigma(i)|\leq K,$
@@
Let $\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{2}$ be classes of boolean functions within $\{0,1\}^{s_{0}}\to\mathbb{R},\{0,1\}^{s_{1}}\to\mathbb{R},\{0,1\}^{s_{2}}\to\mathbb{R}$ respectively. Let $(\mathsf{R}^{(1)},\sigma^{(1)})$ be an explicit $(d_{1},K_{1},\varepsilon_{1})$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{1}$ and $(\mathsf{R}^{(2)},\sigma^{(2)})$ be an explicit $(d_{2},K_{2},\varepsilon_{2})$-weighted pseudorandom reduction from $\mathcal{F}_{1}$ to $\mathcal{F}_{2}$. Then the composition $(\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)},\sigma^{(1)}\cdot\sigma^{(2)}):$
@@
$\sigma^{(1)}\cdot\sigma^{(2)}(i_{1},i_{2}):=\sigma^{(1)}(i_{1})\cdot\sigma^{(2)}(i_{2}),$
@@
Let $\mathsf{R}=\mathsf{R}^{(1)}\circ\mathsf{R}^{(2)}$ and $\sigma=\sigma^{(1)}\cdot\sigma^{(2)}$. For any $f\in\mathcal{F}_{0}$, we have:
$\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d_{1}+d_{2}}}\sum_{i_{1}\in\{0,1\}^{d_{1}},i_{2}\in\{0,1\}^{d_{2}}}\sigma(i_{1},i_{2})\mathbb{E}_{U_{s_{2}}}\left[f(\mathsf{R}_{(i_{1},i_{2})}(U_{s_{2}}))\right]\right|$
$\leq\left|\mathbb{E}_{U}f(U)-\frac{1}{2^{d_{1}}}\sum_{i_{1}\in\{0,1\}^{d_{1}}}\sigma^{(1)}(i_{1})\mathbb{E}_{U_{s_{1}}}\left[f(\mathsf{R}^{(1)}_{i_{1}}(U_{s_{1}}))\right]\right|$
$+\frac{1}{2^{d_{1}}}\sum_{i_{1}\in\{0,1\}^{d_{1}}}|\sigma^{(1)}(i_{1})|\left|\mathbb{E}_{U_{s_{1}}}\left[f(\mathsf{R}^{(1)}_{i_{1}}(U_{s_{1}}))\right]-\frac{1}{2^{d_{2}}}\sum_{i_{2}\in\{0,1\}^{d_{2}}}\sigma^{(2)}(i_{2})\mathbb{E}_{U_{s_{2}}}\left[f(\mathsf{R}^{(2)}_{i_{2}}(U_{s_{2}}))\right]\right|$
@@
Furthermore, $|\sigma(i_{1},i_{2})|\leq|\sigma^{(1)}(i_{1})|\cdot|\sigma^{(2)}(i_{2})|\leq K_{1}K_{2}$. For any $i_{1},i_{2}$, $x\mapsto f(\mathsf{R}^{(1)}_{i_{1}}(x))\in\mathcal{F}_{1}$, sp we can apply $\mathsf{R}^{(2)}_{i_{2}}$ to this mapping, therefore $x\mapsto f(\mathsf{R}_{(i_{1},i_{2})}(x))\in\mathcal{F}_{2}$. The seed length is $d_{1}+d_{2}$. The function is explicit since both components are explicit. ∎
@@
For any positive integer $k$, let $\mathcal{F}_{0},\mathcal{F}_{1},\ldots,\mathcal{F}_{k}$ be classes of boolean functions within $\{0,1\}^{s_{0}}\to\mathbb{R},\{0,1\}^{s_{1}}\to\mathbb{R},\ldots,\{0,1\}^{s_{k}}\to\mathbb{R}$ respectively. Let $(\mathsf{R}^{(1)},\sigma^{(1)}),\ldots,(\mathsf{R}^{(k)},\sigma^{(k)})$ be explicit $(d_{1},K_{1},\varepsilon_{1}),\ldots,(d_{k},K_{k},\varepsilon_{k})$-weighted pseudorandom reductions from $\mathcal{F}_{0}$ to $\mathcal{F}_{1},\ldots,\mathcal{F}_{k}$ respectively. Then the composition $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k)},\sigma^{(1)}\cdot\ldots\cdot\sigma^{(k)}):$
@@
$\sigma^{(1)}\cdot\ldots\cdot\sigma^{(k)}(i_{1},\ldots,i_{k})=\sigma^{(1)}(i_{1})\cdot\ldots\cdot\sigma^{(k)}(i_{k}),$
@@
We prove this by induction on $k$. The base case $k=2$ is shown by Lemma 2.2. Suppose the statement holds for $k-1$, then $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k-1)},\sigma^{(1)}\cdot\ldots\cdot\sigma^{(k-1)})$ is an explicit $(\sum_{i=1}^{k-1}d_{i},\prod_{i=1}^{k-1}K_{i},\sum_{i=1}^{k-1}\left(\prod_{j=1}^{i-1}K_{j}\right)\varepsilon_{i})$-weighted pseudorandom reduction from $\mathcal{F}_{0}$ to $\mathcal{F}_{k-1}$. Then we can apply Lemma 2.2 on $(\mathsf{R}^{(1)}\circ\ldots\circ\mathsf{R}^{(k-1)},\sigma^{(1)}\cdot\ldots\cdot\sigma^{(k-1)})$ and $(\mathsf{R}^{(k)},\sigma^{(k)})$, which gives the desired result. ∎